7. Find the sum of first 22 terms of an AP in which d = 7 and 22nd term is 149.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given, Common difference, d = 7 22nd term, a22 = 149 Sum of first 22 term, S22 = ? By the formula of nth term, an = a+(n−1)d a22 = a+(22−1)d 149 = a+21×7 149 = a+147 a = 2 = First term Sum of n terms, Sn = n/2(a+an) S22 = 22/2 (2+149) = 11×151 = 1661